Find the slope of the straight line x-2y=4.
Answer by tutorcecilia(2144)

x-2y=4 [re-write in the slope-intercept formula (y=mx+b)]
x-x-2y=-x+4
-2y=-x+4
-2y/2=-x/-2+4/-2
y=1/2x-2
Slope = (1/2)